OpenStudy (anonymous):

I DONT KNOW WHY I KEEP GETTING -6 WHEN THE ANSWER FOR THIS PROBLEM IS SUPPOSED TO BE 6!!!! 6-6x(9-9)/1/(2x6)!!! WHEN I PUT IT INTO A Calculator I GET 6 but when i do it out i get -6 PLEASE HELP!!

OpenStudy (anonymous):

lolol calm ndoewn

OpenStudy (anonymous):

sorry, im mad

OpenStudy (anonymous):

well its 6

OpenStudy (anonymous):

lol ths ok. but i say go with the calculator its basicly smarter than allof us:P lol

OpenStudy (anonymous):

9-9=0 therefore 0 divided by anything is =0 there fore 6 remains and thats the answer

OpenStudy (anonymous):

6-6*(9-9)/1/6*2 = 6 - 6*0/1/6*2 = 6-0 =6 ;) i like your hair @Shames ;)

OpenStudy (anonymous):

I dont get how you got 6-0 =6

OpenStudy (anonymous):

And Thanks

OpenStudy (anonymous):

well 9-9 =0 and every thing that is multiplied by 0 is zero and if 0 divided by anything is =0 therefore basically except 6 rest of the stuff is 0 and then 6-0 =0

OpenStudy (anonymous):

oops 6-0 =6 lol

OpenStudy (anonymous):

Thanks @itsmylife

OpenStudy (anonymous):

your welcome